OpenStudy (anonymous):

A rectangle has an area of 48 square feet and a perimeter of 32 feet. Which of the following could be its dimensions?

OpenStudy (anonymous):

area=l*w perimeter=2(l+w) l*w=48 2(l+w)=32 l+w=32/2 l+w=16 l=16-w now substitute l in (l*w=48) (16-w)*w=48 16w-w^2=48 w^2-16w+48=0 w^2-12w-4w+48=0 w(w-12)-4(w-12)=0 (w-12)(w-4)=0 w=12,w=4 so dimensions are 12 and 4